Suppose that the time (in hours) required to repair a machine is an exponentially distributed random variable with parameter λ=0.8, i.e., mean = 1/lambda. What is

(a) the probability that a repair takes less than 77 hours?
Answer #1
here for exponential distribution parameter β = 1/0.8 =1.25
F(x)=P(X<x)=1-e-x/β

probability that a repair takes less than 7 hour :

P(X<7)=1-exp(-7/1.25)= 0.9963